OpenStudy (tylermckinney16):

Say that I have a laser beam that I want to point at a window and I know the sides of the triangle, so I could use trig to solve for that angle.

OpenStudy (tylermckinney16):

Pictures here.

OpenStudy (tylermckinney16):

@Aveline @Directrix @Zarkon @jim_thompson5910

OpenStudy (aveline):

The hypotenuse is \[\sqrt{85}\] So, \[\sin(\theta)=\frac{ 6 }{ \sqrt{85} }\]

OpenStudy (tylermckinney16):

Is that the answer?

OpenStudy (tylermckinney16):

@Aveline

OpenStudy (aveline):

Yes